OpenStudy (anonymous):

PLEASE HELP expand the logarithmic expression (will input expression w/ the equation creator)

OpenStudy (anonymous):

\[\log _{6} c/4\]

OpenStudy (anonymous):

Do you want the series expansion of log or do you want to write it as \[\log_{6} c- \log_{6} 4\] (log a/b)=log a- log b

OpenStudy (anonymous):

2nd, to give me a more simple answer, not series expansion

OpenStudy (anonymous):

Mth3v4 I cant see what you wrote can post a link please is says canot conect to web

OpenStudy (anonymous):

So just use log(a/b)=log a- log b the answer i have already mentioned.

OpenStudy (anonymous):

so what the expanded form be log base 6 c over log base 6 4?

OpenStudy (anonymous):

What you have posted is log base 6 c over 4 NOT log base 6 c over log base 6 4.

OpenStudy (anonymous):

so i have it flipped around, the denominator i have is the numerator and vise-versa?

OpenStudy (anonymous):

What on earth are you saying? The answer is just \[\log_{6}c-\log_{6}4 \]

OpenStudy (anonymous):

haha sorry, i just dont see that in mmy multiple choice answers, what is it in series?

OpenStudy (anonymous):

im sorry, thanks for helping me

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

nvm i got it, thanks!!